Calculate the probability for the following situation, then select the correct answer:
You are tossing a coin, then rolling a die, then drawing a card from a deck of cards. What is the probability that you will get: a head AND an odd number on the die AND a card greater than 3 (assume the ace is equal to 1) from the deck?

Answer by greenestamps(13195)



P(head) = 1/2
P(odd) = 3/6 = 1/2
P(greater than 3) = 40/52 = 10/13

Since all three are required to happen, the overall probability is the product of those three probabilities.